### v3

# Time-Constrained Erasure Correction for Data Recovery in UAV-LoRa-WuR Networks

Abstract
We consider a LoRa sensor network featuring a UAV-mounted gateway for collecting sensor data and wake-up radios (WuR) for activating the sensors upon the UAV’s arrival. To recover lost data on the sensor-to-UAV links, we design two erasure-correction mechanisms based on fountain coding and message replication. The design process accounts for the time constraints imposed by the UAV’s limited hovering duration.
Index Terms:
LoRa, UAV, Wake-Up Radio, Data Recovery
I Introduction
The LoRa technology suites IoT applications that require long-distance transmission of sensor data. By adjusting its spreading factor (SF), LoRa can provide large communication ranges at the cost of data rate and energy [1]. To reduce the energy expenditure when the receiver is far away, and to enable communications when it is not reachable, UAV-aided LoRa has been explored [2]–[5]. The idea is to employ a UAV-mounted LoRa gateway that hovers over a cluster of sensors from time to time, so that the sensors can transmit their data to the UAV over a shorter distance. The UAV relays the sensor data over a high-capacity link to the final destination.
Communication protocols for such settings must address questions such as how to alert the sensors when the UAV arrives so that they can transmit their data, how the sensors should access the shared uplink channel within the UAV’s limited hovering time, and how to ensure acceptable reliability levels on the uplink. Among related works, [2]–[4] do not consider the issue of alerting the sensors. It is addressed in [5], where the sensors and the UAV are equipped with WuR modules in addition to the main radio (i.e., the LoRa transceiver). A WuR consumes negligible power while scanning for and receiving signals [6]. When the UAV is not present, the sensors keeps their main radio switched off, while keeping the WuR on. Upon arrival, the UAV transmits a series of wake-up beacons (WUBs) from its WuR module to awaken and synchronize the sensors, which then power on their main radio and access the shared channel with random selection of transmission slots, frequency band, and SF.
The simplicity of the method in [5] makes it well-suited for deployment using unsophisticated sensor nodes. However, as with all random access schemes, the transmissions are prone to collisions and losses. As a remedy, we devise two erasure-correction schemes for improving message delivery. One scheme employs fountain coding, and transmits redundancy in the form of random linear combinations of the sensor readings. The other scheme performs message replication by transmitting multiple copies of a reading. Unlike conventional redundancy transmission schemes, the limited hovering time of the UAV restricts the transmission capacity of a sensor. Another constraint arises from the sensor’s limited energy budget. These constraints are considered in our design.
II System Model
We consider a cluster of $n$ sensors or end devices (EDs), each containing a WuR and a LoRa transceiver and performing a reading every $T_{e}$ hours. The UAV visits the cluster every $T_{u}$ hours. Thus, every ED has $m\!=\!T_{u}/T_{e}$ readings (messages) to send to the UAV. To wake up the sensors, the UAV transmits a sequence of WUBs, each having a reception probability $P_{b}$. Upon waking, a sensor transmits its data to the UAV in a LoRa frame using one of $N_{f}$ frequency bands.
LoRa frame characteristics depend on the SF, which is an integer between $7$ and $12$. The higher the SF, the greater the immunity against thermal noise and longer the frame duration. The SFs are quasi-orthogonal; that is, a frame is highly likely to survive interference from a colliding frame having a different SF, but is highly likely to be lost upon collision with the same SF. For brevity, here we assume as in [5] that different-SF collisions do not enforce losses, whereas same-SF collisions always do. This is a reasonable assumption under mild fading and comparable distances from all sensors to the UAV [5]. More precise models for LoRa’s collision behavior, such as the one given in [7]), can be easily incorporated into our analysis – but at the expense of lengthening the derivations and exceeding this letter’s page limits.
Sensor Wake-Up and Uplink Channel Access
The method of [5] is employed for this purpose. It involves slotted communications from the EDs to the UAV with slot length $T_{s}$. The UAV hovers over the EDs for $N_{s}$ slots, numbered 0 through $N_{s}\!-\!1$. At every slot beginning, the UAV broadcasts a WUB whose data field contains the current slot number $i$ and the value of $N_{s}$. Once awake, an ED transmits its messages to the UAV using one of the methods to be described in Section III. For every transmission on the ED-to-UAV uplink, one of the $N_{f}$ frequency bands is chosen uniformly at random. Likewise, the SF is chosen uniformly at random from the set $\mathbf{K}\!=\!\{7,8,\ldots,K_{m}\}$, where $127\!\leq\!K_{m}\!\leq\!$ 12. Thus, the probability of selecting SF $k\in\mathbf{K}$ is $\eta_{k}=1/(K_{m}-6).$
Sensor Transmission Limits
The energy budget of a sensor imposes a limit on how many frames the sensor can send per data-collection cycle. Let $L_{f}(k)$ be the length of a LoRa frame using SF $k$, and let $P_{t}$ be a sensor’s power consumption while transmitting. Thus, the average energy consumption per frame is $P_{t}\overline{L}_{f}$, where $\overline{L}_{f}$ is the average frame length, given by $\overline{L}_{f}=\sum_{k=7}^{K_{m}}\eta_{k}L_{f}(k)$ The maximum number of frames a sensor can send per data-collection cycle is $N_{\max}={E_{\max}}/{(P_{t}\overline{L}_{f})},$ where $E_{\max}$ is the maximum allowed energy expenditure per cycle. It is a design parameter that depends on battery capacity and expected sensor lifetime. In the following, we assume that $N_{\max}\!>\!m$.
Fountain Coding
We consider a random linear fountain code [9]. Let $\psi_{0},\psi_{1},\ldots,\psi_{m-1}$} denote an ED’s source messages. Fountain encoding maps them to $m+\varepsilon$ coded messages {$c_{0},c_{1},\ldots,c_{m+\varepsilon-1}$}, where $c_{j}=\sum_{l=0}^{m-1}\alpha_{l}^{(j)}\psi_{l},$ where the coefficients $\alpha_{l}^{(j)}$ are chosen at random from a Galois field of a given order $q$ (GF($q$)). The original $m$ messages can be decoded from a set of $z$ coded messages if the set contains $m$ linearly independent combinations of the messages. This occurs with probability $P_{\mathrm{dec}}(z)\!=\!\prod_{v=0}^{m-1}(1-q^{v-z})$ for $z\!\geq\!m$, and 0 for $z\!<\!m$ [9].
III Redundancy Transmission Schemes
Our proposed schemes employ the transmission of up to $m\!+\!\varepsilon$ frames per cycle, where the design parameter $\varepsilon\!\leq\!N_{\max}\!-\!m$ represents the amount of redundancy. We define $N(i)\!=\!N_{s}\!-\!i$ and $\gamma(i)\!=\!N(i)\!-\!m.$ Thus, for an ED that wakes up in slot $i$, there are $N(i)$ slots remaining in the session, of which $\gamma(i)$ are available for redundancy transmission. Note that $\gamma(i)$ may be 0 or negative depending on the value of $i$. The two erasure-correction schemes operate as follows:
III-A Fountain Coding
An ED first checks if there are enough slots available to transmit $\varepsilon$ redundant frames, that is, whether $\gamma(i)\!\geq\!\varepsilon$.
If $\gamma(i)\!\geq\!\varepsilon$: The ED encodes the $m$ source messages into $m\!+\!\varepsilon$ fountain-coded messages and transmits them over $m\!+\!\varepsilon$ randomly chosen slots out of the $N(i)$ available slots.
If $\gamma(i)\!<\!\varepsilon$: Fountain coding is not applied. Instead, if $m\!\leq\!N(i)$, the $m$ messages are transmitted over $m$ randomly chosen slots. If $m\!>\!N(i)$, then the ED randomly selects $N(i)$ messages and sends them over the $N(i)$ slots. Messages not selected are discarded and deemed lost.
III-B Message Replication
This scheme makes transmission decisions as follows:
$\gamma(i)\!\geq\!0$: Let $\hat{\varepsilon}(i)\!=\!\min\{\gamma(i),\varepsilon\},$ and $m_{q}$ and $m_{r}$ be the quotient and the remainder, respectively, when $\hat{\varepsilon}(i)$ is divided by $m$. Each of the $m$ messages is sent $m_{q}$ times; further, $m_{r}$ messages are randomly picked for one additional transmission.
$\gamma(i)\!<\!0$: This implies $N(i)\!<\!m$. The ED randomly selects $N(i)$ messages and sends them over the $N(i)$ slots. Messages not selected are discarded and deemed lost.
IV Performance Analysis
Let the message delivery probability denote the probability that an arbitrary message from an ED is delivered to the control station. We derive it below for the two schemes.
IV-A Fountain Coding
First note that the probability mass function for the slot in which an ED wakes up (denoted by $W$) is
$P_{W}(i)=(1-P_{b})^{i}P_{b}$ (1)
for $10\leq i\leq N_{s}-$ 1, and 0 otherwise.
Lemma 1.
An ED that wakes up in slot $i$ transmits a frame to the UAV in slot $s$ with probability
$P_{\mathrm{tx}}(s|W=i)=\begin{cases}\frac{m+\varepsilon}{N(i)},\quad&s\geq i,\gamma(i)\!\geq\!\varepsilon\\ \min\left\{\frac{m}{N(i)},1\right\},\quad&s\geq i,\gamma(i)\!<\!\varepsilon\\ 0\quad&s<i.\end{cases}$ (2)
Proof:
By definition, the ED has $N(i)$ slots available for use. If $\gamma(i)\!\geq\!\varepsilon$, then $m\!+\!\varepsilon$ coded messages are randomly distributed over $N(i)$ slots; hence the probability of choosing a certain slot is $(m\!+\!\varepsilon)/N(i)$. If $\gamma(i)\!<\!\varepsilon$, then $m$ uncoded messages are to be distributed over $N(i)$ slots; the probability of choosing a certain slot is either $m/N(i)$ (if $m\!\leq\!N(i)$), or 1 (if $m\!>\!N(i)$, meaning all slots are used and some messages are dropped.) This is succinctly expressed as $\min\{m/N(i),1\}$. Of course, the probability of choosing slot $s$ is 0 if $s\!<\!i$. ∎
Let the transmission success probability (denoted $\zeta(s)$) be the probability that a frame sent in slot $s$ is successfully received at the UAV. Consider such a frame $\mathrm{F}$ from an arbitrary ED $\mathrm{E}$. Then $\zeta(s)$ is the probability that no other ED transmits in slot $s$ using the same SF and the same frequency band. It can be determined as follows:
Lemma 2.
The probability that a given ED $\mathrm{E}^{\prime}$ transmits a frame in a given slot $s$ is
$P_{\mathrm{col}}(s)$ $=\sum_{j=0}^{\min\{N_{s}-m-\varepsilon,s\}}\frac{m+\varepsilon}{N(j)}P_{W}(j)$
$+\theta_{s}\sum_{j=N_{s}-m-\varepsilon+1}^{s}\min\left\{\frac{m}{N(j)},1\right\}P_{W}(j),$ (3)
where
$\theta_{s}=\begin{cases}1,\quad&s>N_{s}-m-\varepsilon\\ 0,\quad&\text{otherwise}.\end{cases}$ (4)
Proof:
Suppose $\mathrm{E}^{\prime}$ wakes up in slot $j$. Then $\gamma(j)\!\geq\!\varepsilon$ if $j\!\leq\!N_{s}\!-\!m\!-\!\varepsilon$. From the argument in Lemma 1, the probability of selecting slot $s$ in this case is $(m\!+\!\varepsilon)/N(j)$. The first term in (2) averages this probability over all $j$ such that $\gamma(j)\!\geq\!\varepsilon$. The $\min$ operation in the upper limit of the summation ensures that we consider only the values of $j$ up to $s$ (since an ED waking up after slot $s$ cannot transmit in slot $s$.) Similarly, the second summation accounts for the transmission probabilities for all $j$ such that $\gamma(j)\!<\!\varepsilon$. The multiplier $\theta_{s}$ ensures that we consider the second summation only when there exists a $j$ that simultaneously satisfies both $\gamma(j)\!<\!\varepsilon$ and $j\!\leq\!s$. ∎
The transmission from $\mathrm{E}^{\prime}$ must employ the same SF and the same channel as $\mathrm{F}$ to cause a loss. Note that these occur with probability $\eta_{k}$ and $1/N_{f}$, respectively. Hence, the probability that one given ED (namely, $\mathrm{E}^{\prime}$) enforces the loss of $\mathrm{F}$ is
$\mathcal{L}(s)=\eta_{k}P_{\mathrm{col}}(s)/N_{f}.$ (5)
Since there are $n\!-\!1$ EDs excluding the sender of $\mathrm{F}$, the transmission success probability for $\mathrm{F}$ is given by
$\zeta(s)=(1-\mathcal{L}(s))^{n-1}=\left(1-\eta_{k}P_{\mathrm{col}}(s)/N_{f}\right)^{n-1}.$ (6)
To derive the message delivery probability, consider an arbitrary message $\psi$ from an arbitrary ED $\mathrm{E}$. Suppose $\mathrm{E}$ wakes up in slot $i$. Then the following events can occur:
IV-A1 $\gamma(i)\!\geq\!\varepsilon$
In this case, $\mathrm{E}$ performs fountain coding and transmits $m\!+\!\varepsilon$ coded messages. Thus, the MDP for $\psi$ is the probability that the received set of coded messages is decodable. The message delivery probability then becomes
$\mathcal{S}_{1}(i)\!=\!\sum_{z=m}^{m+\varepsilon}\xi(z,i)P_{\mathrm{dec}}(z)\!=\!\sum_{z=m}^{m+\varepsilon}\xi(z,i)\prod_{v=0}^{m-1}(1\!-\!q^{v-z}),$ (7)
where $\xi(z,i)$ is the probability that $z$ out of the $m\!+\!\varepsilon$ transmissions are successful. Finding an exact expression for $\xi(z,i)$ is complicated due to the slot-dependent nature of the transmission success probability $\zeta(s)$. However, a fairly accurate approximation may be obtained by assuming that the transmission success events over different slots are mutually independent and has a fixed probability $\hat{\zeta}$, and then using the binomial mass function
$\xi(z,i)\approx{m+\varepsilon\choose z}\hat{\zeta}^{z}(1-\hat{\zeta})^{m+\varepsilon-z}.$ (8)
The parameter $\hat{\zeta}$ is obtained by averaging the transmission success probabilities for the $N(i)$ slots that are available to the ED. That is
$\hat{\zeta}=\frac{1}{N(i)}\sum_{s=i}^{N_{s}-1}\zeta(s).$ (9)
IV-A2 $\gamma(i)\!<\!\varepsilon$
In this case, fountain coding is not performed, and the message $\psi$ is sent as it is. The probability that $\psi$ is sent over slot $s$ is given by [5]
$\mathcal{T}(s,i)=\frac{1}{N(i)}\min\left\{\frac{N(i)}{m},1\right\},\quad i\leq s\leq N_{s}-1.$ (10)
The message delivery probability is thus
$\mathcal{S}_{2}(i)$ $=\sum_{s=i}^{N_{s}-1}\mathcal{T}(s,i)\zeta(s).$ (11)
The overall message delivery probability is
$\mathcal{S}=\sum_{i=0}^{N_{s}-m-\varepsilon}P_{W}(i)\mathcal{S}_{1}(i)+\sum_{i=N_{s}-m-\varepsilon+1}^{N_{s}-1}P_{W}(i)\mathcal{S}_{2}(i),$ (12)
where the first summation is over those values of the wake-up instant $i$ such that $\gamma(i)\!\geq\!\varepsilon$, and the second one is for $\gamma(i)\!<\!\varepsilon$.
IV-B Message Replication
Lemma 3.
An ED that wakes up in slot $i$ transmits a frame to the UAV in slot $s$ with probability
$\hat{P}_{\mathrm{tx}}(s|W=i)=\begin{cases}\frac{m+\hat{\varepsilon}(i)}{N(i)},\quad&s\geq i,\gamma(i)>0\\ \min\left\{\frac{m}{N(i)},1\right\},\quad&s\geq i,\gamma(i)\leq 0\\ 0\quad&s<i.\end{cases}$ (13)
Proof:
The proof is same as that of Lemma 1, except that an ED transmits redundancy for any $\gamma(i)\!>\!0$, and $\hat{\varepsilon}(i)$ is the number of redundant frames sent. ∎
Lemma 4.
The probability that a given ED $\mathrm{E}^{\prime}$ transmits a frame in a given slot $s$ is
$\hat{P}_{\mathrm{col}}(s)$ $=\sum_{j=0}^{\min\{N_{s}-m,s\}}\frac{m+\hat{\varepsilon}(j)}{N(j)}P_{W}(j)$
$+\hat{\theta}_{s}\sum_{j=N_{s}-m+1}^{s}\min\left\{\frac{m}{N(j)},1\right\}P_{W}(j),$ (14)
where
$\hat{\theta}_{s}=\begin{cases}1,\quad&s>N_{s}-m\\ 0,\quad&\text{otherwise}.\end{cases}$ (15)
Proof:
The proof is same as that of Lemma 2, except that an ED transmits redundancy for any $\gamma(i)\!>\!0$, and $\hat{\varepsilon}(i)$ is the number of redundant frames sent. ∎
Using the same argument as in Section IV-A, the transmission success probability for a frame is expressed as
$\hat{\zeta}(s)=\left(1-\eta_{k}\hat{P}_{\mathrm{col}}(s)/N_{f}\right)^{n-1}.$ (16)
To derive the message delivery probability, consider an arbitrary message $\psi$ from an arbitrary ED $\mathrm{E}$. Suppose $\mathrm{E}$ wakes up in a slot $i$. Then the following events can occur:
IV-B1 $\gamma(i)>0$
Of the $m$ messages, $m_{r}(i)$ are sent $m_{q}(i)\!+\!1$ times, and the remaining $m\!-\!m_{r}(i)$ are sent $m_{q}(i)$ times. Thus, $\mathcal{P}_{1}(i)\!=\!(m\!-\!m_{r}(i))/m$ and $\mathcal{P}_{2}(i)\!=\!m_{r}(i)/m$ denote the probability that a given message is sent $m_{q}(i)$ times and $m_{q}(i)\!+\!1$ times, respectively. The corresponding probabilities that at least one copy of the message is delivered are approximately $1-(1-\hat{\zeta})^{m_{q}(i)}$ and $1-(1-\hat{\zeta})^{m_{q}(i)\!+\!1}$, where $\hat{\zeta}$ is defined as in (9).
$\hat{\mathcal{S}}_{1}(i)=$ $\mathcal{P}_{1}(i)\left(1-(1-\hat{\zeta})^{m_{q}(i)}\right)$
$+\mathcal{P}_{2}(i)\left(1-(1-\hat{\zeta})^{m_{q}(i)+1}\right).$ (17)
IV-B2 $\gamma(i)\leq 0$
In this case, the probability that a given message is sent over slot $s$, and the subsequent message delivery probability are the same as in the fountain-coding scheme, and are given by (10) and (11), respectively.
Using a similar argument as used for (12), the overall message delivery probability for message replication is
$\hat{\mathcal{S}}=\sum_{i=0}^{N_{s}-m}P_{W}(i)\hat{\mathcal{S}}_{1}(i)+\sum_{i=N_{s}-m+1}^{N_{s}-1}P_{W}(i)\hat{\mathcal{S}}_{2}(i),$ (18)
where $\hat{\mathcal{S}}_{2}(i)$ is the same as $\mathcal{S}_{2}(i)$ given by (11).
V Numerical Results
Figure 1: Reliability vs. UAV’s hovering time $(\varepsilon\!=\!4$). Circles show simulated values, lines show analytical results.
We compare the proposed schemes with the uncoded scheme of [5]. Unless otherwise specified, $n\!=\!20$, $T_{e}\!=\!1$ hour, $T_{u}\!=\!5$ hours (hence $m\!=\!5$), $N_{f}\!=\!8$, $P_{b}\!=\!0.25$, $\mathbf{K}=\{7,8,9\}$, and $q\!=\!256$. We keep $P_{b}$ small to model an unfavorable scenario where WUB receptions face significant challenge, possibly owing to the UAV’s hovering altitude.
Fig. 1 shows the message delivery probabilities for $\varepsilon\!=\!4$. We observe that when the UAV hovers for 15 or more slots, the proposed schemes noticeably outperform the baseline. The relative improvement is maximum for $N_{s}\!\approx\!30$. For $N_{s}\!>\!18$, fountain coding outperforms message replication. As $N_{s}\!\rightarrow\!\infty$, the three schemes are expected to asymptotically converge, since the collision probability becomes arbitrarily small and all messages are successfully delivered. However, for smaller (and more realistic) hovering times, redundancy is beneficial.
The impact of $\varepsilon$ is further examined in Fig. 2, where results for $\varepsilon\!=\!1$ and $\varepsilon\!=\!3$ are shown. For $\varepsilon\!=\!3$, the same conclusions as made from Fig. 1 are applicable. However, for $\varepsilon\!=\!1$, message replication provides only marginal improvement over the baseline. Further, fountain coding suffers significant performance deterioration relative to the other two schemes when $N_{s}\!<\!70$. This is attributed to the fact that when $\epsilon\!=\!1$, the set of coded messages is not decodable whenever there is more than a single erasure. Even with a single or no erasure, decoding may fail due to the rank of the set of received packets being less than $m$; however, with a large $q$ (e.g., as in our case), that probability is quite small. For $N_{s}\!>\!70$, frame losses become rarer, and fountain coding outperforms both message replication as well as the baseline scheme for $\varepsilon\!=\!1$.
Fig. 3 shows the impact of the number of EDs in the cluster. With increasing node density, more collisions occur, and the performance of each scheme degrades. But with $\varepsilon\!=\!3$, both schemes outperform the baseline, with fountain coding being significantly better. For $\varepsilon\!=\!1$, message replication performs virtually the same as the baseline, whereas fountain coding is worse than the baseline for high node densities (due to more frame losses leading to rank deficiency) and is better when there are fewer EDs.
Figure 2: Reliability as a function of UAV hovering time.
Figure 3: Reliability as a function of node density $n$ ($N_{s}\!=\!60$.)
VI Conclusion
We described two erasure-correction schemes for data recovery in UAV-LoRa-WuR networks. Our results show that unless the maximum number for redundant frames a sensor can send per data-collection cycle is very small, erasure coding provides noticeable improvements over an uncoded transmissions. Whether to employ coding – and if so, which type – should be determined based on the sensors’ energy budget (which dictates the maximum redundancy), the UAV’s hovering time, and the node density. The analytical framework presented above aids in this decision making.